OpenStudy (anonymous):

Write the equation of the line perpendicular to y = 2x + 1 and passing through the point (-2,3).

OpenStudy (mrnood):

ok the general form of a straight line is y=mx + c compare that to YOUR equation y= 2x+1 what is the value equivalent to m, in YOUR equation?

OpenStudy (anonymous):

the value that is equal to m is 2 . but i need perpendicular

OpenStudy (mrnood):

OK - we're getting there m is the slope of the line the slope of a line perpendicular to that is -1/m so if the slope is 2 - what is the slope of the perpendicular line?

OpenStudy (anonymous):

-1/2

OpenStudy (mrnood):

ok good so the equation of the perpendicular line is of the form y = -1/2 x +b but you are told that when x=-2 then y =3 (because the point given lies on the line) so put those values into the equation, and you can then work out what b is

OpenStudy (anonymous):

3=2(-2)+1

OpenStudy (mrnood):

no - it is y= (-1/2) x +b

OpenStudy (mrnood):

put y = 3, and x= -2 you will then have an equation with b as the unknown value

OpenStudy (anonymous):

b=2

OpenStudy (mrnood):

so you now know m and b so write the new equation y = mx+b

OpenStudy (anonymous):

y=-1/2+2

OpenStudy (mrnood):

yes - just about... it's y = -1/2 x + 2 well done. Just remember - if you have a line of slope m then the perpendicular line has slope -1/m